OpenStudy (anonymous):

Which is shown in increasing order? zero to the twelfth power, 2 to the third power, 4 to the second power, 9 to the zero power, ten to the fourth power, twelve to the second power 9 to the zero power, 4 to the second power, twelve to the second power, ten to the fourth power, twelve to the twelfth power 9 to the zero power, zero to the twelfth power, 2 to the third power, 4 to the second power, ten to the fourth power, twelve to the second power zero to the twelfth power, 9 to the zero power, 2 to the third power, 4 to the second power, twelve to the second power, ten to the fourt

OpenStudy (blockcolder):

0^12=0 2^3=8 4^2=16 9^0=1 10^4=10,000 12^2=144 Now, arrange these in increasing order.

OpenStudy (anonymous):

Yes arranging them in increasing order, it would be A.

OpenStudy (anonymous):

10,000,144,16,8,1,0

OpenStudy (anonymous):

zero to the twelfth power, 2 to the third power, 4 to the second power, 9 to the zero power, ten to the fourth power, twelve to the second power... In math terms... 0-6-8-9-40-144

OpenStudy (anonymous):

And all those numbers are in ascending/increasing order:)

OpenStudy (anonymous):

is it a

OpenStudy (blockcolder):

@Rain4me I think you mistook "power" for multiplication. :) @kiddpolo43 It's D.

OpenStudy (anonymous):

Yes it is A...but also understand why so Im not just telling you the answer..

OpenStudy (anonymous):

Oh your right..I did in parts...

OpenStudy (anonymous):

So it is D.

OpenStudy (anonymous):

okay is it a or d

OpenStudy (anonymous):

I would agree with blockcolder and say D. It the powers..not multiplication like I was doing..

OpenStudy (anonymous):

okay that what i thought it was